# Bayesian competing risks survival modeling for assessing the cause of death of patients with heart failure

### v1

@@ Abstract
Competing risk models are survival models with several events of interest acting in competition and whose occurrence is only observed for the event that occurs first in time. This paper presents a Bayesian approach to these models in which the issue of model selection is treated in a special way by proposing generalizations of some of the Bayesian procedures used in univariate survival analysis. This research is motivated by a study on the survival of patients with hearth failure undergoing cardiac resynchronization therapy, a procedure which involves the implant of a device to stabilize the heartbeat. Two different causes of causes of death have been considered: cardiovascular and non-cardiovascular, and a set of baseline covariates are examined in order to better understand their relationship with both causes of death. Model selection procedures and model checking analyses have been implemented and assessed. The posterior distribution of some relevant outputs such as the overall survival function, cumulative incidence functions, and transition probabilities have been computed and discussed.
keywords:
Cardioscular and non-cardiovascular death; Conditional Predictive Ordinate; Model uncertainty; Transition probabilities.
@@ 1 Introduction
The statistical framework for the aforementioned objectives is Survival Analysis. In particular, competing risk models in which the event of interest, in our case death, can occur from a variety of competing causes. The statistical literature is abundant regarding procedures for modeling competing survival scenarios (Putter et al., 2007; Chen et al., 2014). In particular, we will focus on cause-specific risk modelling to quantitatively assess the behaviour of the survival process through the cause-specific hazard function for each cause of death, the overall survival function for the time in which death occurs, and the cumulative incidence function of that time due to each of the causes considered among others (Chen et al., 2014). We also incorporate a look at the modelling from the point of view of multistate models (Andersen et al., 2001). These are stochastic processes of continuous time and discrete state space representing the different health conditions considered. We discuss the transition probabilities between the different states and specially conditional transition probabilities when it is known that a patient remains alive a certain time after the CRT treatment. We will work within the Bayesian statistical methodology. This allows quantifying all sources of uncertainty associated with a statistical problem in a natural way through probability distributions.
The rest of the paper is organized as follows. Section 2 presents a retrospective follow-up cardiac resynchronization therapy study and a brief description of the subsequent dataset. Section 3 contains general information about the competing risk model, the main elements of the cause-specific risk approach, and the general Bayesian inferential process. Subsection 3.3, devoted to model uncertainty, presents a special variable selection scenario. In this topic, we extend the proposal by García García-Donato et al (2023) in Cox regression models for computing Bayes factors to competing risk models and employ spike-and-slab selection methods. In addition, we also expand the definition of conditional predictive ordinate (CPO) (Ibrahim et al., 2001) to the general competing risk framework which is needed in order to apply the logarithm of the pseudomarginal likelihood as a tool for model assessment. Section 4 returns to the study of CRT therapy and applies the procedures and proposals of Section 3 to the data of this problem. Finally, Section 5 includes some conclusions.
@@ 2 The cardiac resynchronization therapy study
Clinical follow-up was completed in all patients, from the time each individual underwent CRT implantation (time zero) until death, cardiovascular (CV) or non-cardiovascular (N-CV), or until the date of the end of the follow-up. Survival times were observed in the first case and were right-censored in the second. The median and interquartile ranges of survival times were 2.75 and (1.2, 5.2) years, respectively.
@@
Table 1: Descriptive summary of study covariates for censored individuals, for those who died, from CV or N-CV causes, and for all of the together. In the case of quantitative covariates, the sample mean and standard deviation are presented; for categorical variables, the number of patients in the sample in each category and their percentage are reported.
@@
Previous AF 83 (40.89%) 31 (56.36%) 24 (63.16%) 138 (46.62%)
@@
The standard Kaplan-Meier procedures for describing survival probabilities do not provide proper results in the context of competing risks. In these settings, an alternative measure is the cumulative incidence function for each cause, defined as the cumulative probability of failure from a specific cause before a certain time. Figure 1 presents the cumulative incidence function for CV and N-CV death computed from the R package cmprsk (Gray, 2022). CV and N-CV death behave quite similarly, although cardiovascular death is always associated with a higher probability. It is important to note that after five years of the CRT intervention the difference between the two probabilities increases: between 0.15 and 0.20 for N-CV death and between 0.20-0.25 for CV death. From 7-8 years onwards there is an abrupt rise in CV death, while CV remains at around 0.2. The small number of patients we have from 10 years onwards does not produce consistent informative signals.
@@ 3 Bayesian competing risks modelling
3.1 Competing risk sampling model
Competing risk models are survival models in which the initiating event of the process contemplate different terminating events, which compete with each other to be the event that eventually occurs. There are different models suitable for analyzing this type of problems. The most popular, and the one we will use in our study, is the cause-specific hazard model. We assume $K$ terminating events or causes of failure.
We define $T_{k}$ as the random variable that describes the time to the terminating event $k$, and $\delta$ as the indicator variable for the event that has occurred. In particular, $\delta=k$ when the event $k$ has occurred and $\delta=0$ indicating the situation in which no event of interest has been recorded during the study period and, therefore, we have right censored information. Let $T=\mbox{min}\{T_{1},\ldots,T_{K}\}$. Cause-specific hazard modelling uses as a basic concept the joint distribution of $T$ and $\delta$ through the so-called cause-specific hazard function, which represents the hazard of failing from a given event in the company of all other competing events. The cause-specific hazard function for $T$ due to cause $k$ is defined as
$h_{k}(t\mid\bm{\theta}_{k})=\lim_{\varDelta t\to 0}\frac{P(t\leq T<t+\varDelta t,\delta=k\mid T\geq t,\,\bm{\theta}_{k})}{\varDelta t},$ (1)
where $\bm{\theta}_{k}$ is the subsequent parametric vector. It is important to remember that our inferential framework is Bayesian statistics. Consequently, the uncertainty associated with the parameters will be expressed in probabilistic terms and for this reason we will express in terms of conditional probability distributions all probabilistic elements, such as $h_{k}(t\mid\bm{\theta_{k}})$, that depend on parameters.
The most common modeling of the cause-specific hazard function for cause $k$ is through a proportional hazards model (Chen et al., 2014) as follows
$h_{k}(t\mid\bm{\theta}_{k})=h_{0,k}(t\mid\bm{\phi}_{k})\>\exp\{\bm{x}^{\top}_{k}\,\bm{\beta}_{k}\},$ (2)
where $h_{0,k}(t\mid\bm{\phi_{k}})$ is a baseline hazard function that depends on time and a parametric vector $\bm{\phi}_{k}$. It expresses the shape of the cause-specific hazard function. The vector $\bm{\beta}_{k}$ represents the regression coefficients associated to covariates $\bm{x}_{k}$ of covariates, and $\bm{\theta}_{k}=(\bm{\phi}_{k},\bm{\beta}_{k})$. Baseline hazard functions can be established through parametric or semiparametric approaches. Parametric models generate functions with rigid shapes that cannot admit irregular behavior. On the other hand, semi-parametric modeling provides very flexible functions but suffer from overfitting and instability (Ibrahim et al., 2014).
@@ 3.1 Competing risk sampling model
where now $\bm{\theta}=(\bm{\theta}_{1},\ldots,\bm{\theta}_{K})$.
@@
$F_{k}(t\mid\bm{\theta})=P(T\leq t,\>\delta=k\mid\bm{\theta})=\int_{0}^{t}h_{k}(u\mid\bm{\theta_{k}})\>S(u\mid\bm{\theta})\>\mbox{d}u.$
This function provides the probability associated with the occurrence of each of the competing events, $P(\delta=k\mid\bm{\theta})=\lim_{t\to\infty}F_{k}(t\mid\bm{\theta})$.
Competing risk models can also be seen as a particular multi-state model. In general, a multi-state process is a continuous time stochastic process $\{Z(t),t\geq 0\}$ with a finite state space $\Omega=\{0,1,\ldots,K\}$, where $Z(t)=k$ indicates that the process is in state $k$ at time $t$, or in terms of survival analysis that event $k$ has occurred at some time instant less than $t$. The initial distribution of the process is $P(Z(0)=1)=1$ and relevant transition probabilities between states are
$p_{(0,0)}(0,t\mid\bm{\theta})$ $=P(Z(t)=0\mid Z(0)=0,\bm{\theta})=P(T>t\mid\bm{\theta})=S(t\mid\bm{\theta}),$
$p_{(0,k)}(0,t\mid\bm{\theta})$ $=P(Z(t)=k\mid Z(0)=0,\bm{\theta})=P(T\leq t,\delta=k\mid\bm{\theta})=F_{k}(t\mid\bm{\theta}),\,\,k\neq 0.$
@@
$p_{(0,0)}(s,t\mid\bm{\theta})=$ $\,\frac{P(Z(t)=0,Z(s)=0\mid\bm{\theta})}{P(Z(s)=0\mid\bm{\theta})}=\frac{S(t\mid\bm{\theta})}{S(s\mid\bm{\theta})},\,t\geq s,$
$p_{(0,k)}(s,t\mid\theta)=$ $\,\frac{P(Z(t)=k,Z(s)=0\mid\bm{\theta})}{P(Z(s)=0\mid\bm{\theta})}=\frac{P(s<T<t,\delta=k\mid\bm{\theta})}{S(s\mid\bm{\theta})}$
$=$ $\,\frac{F_{k}(t\mid\bm{\theta})-F_{k}(s\mid\bm{\theta})}{S(s\mid\bm{\theta})},\,t\geq s,\,k\neq 0.$
@@ 3.2 Prior and posterior distributions
Bayesian models include a sampling model and a prior distribution $\pi(\bm{\theta})$ for the unknown parameters that describes how much, little or no information is available about $\bm{\theta}$. We will work in a general scenario of prior independence between the parameters associated with the different competing risks as well as between the parameters governing the time-dependent structure of each hazard function and the regression coefficients of the variables, $\pi(\bm{\theta})=\prod_{k=1}^{K}\,\pi(\bm{\phi}_{k},\bm{\beta}_{k})=\prod_{k=1}^{K}\,\pi(\bm{\phi}_{k})\,\pi(\bm{\beta}_{k})$.
The database is built on the basis that survival times due to one of the causes provides censored information on the survival times of the other competing causes. The information provided by the data $\mathcal{D}$ is described in terms of the likelihood function of $\bm{\theta}$, $L(\bm{\theta})$, which is calculated through the hazard function (Chen et al., 2014). The prior distribution and the likelihood function are combined to obtain the posterior distribution of $\bm{\theta}$, $\pi(\bm{\theta}\mid\mathcal{D})$. This distribution is not analytical but an approximate sample $\{\bm{\theta}^{(1)},\ldots,\bm{\theta}^{(M)}\}$ of such a distribution can be obtained through Markov chain Monte Carlo, MCMC, methods. This distribution includes all available information and is the starting point for the analysis of the problem.
The overall survival function, cumulative incidence function, probability associated with the occurrence of each competing event as well as transitions probabilities presented before are relevant outputs that depend on $\bm{\theta}$. Consequently, we can use the posterior approximate sample from $\pi(\bm{\theta}\mid\mathcal{D})$ to generate approximate random samples for their subsequent posteriors, $\pi(S(t\mid\bm{\theta})\mid\mathcal{D})$, $\pi(F_{k}(t\mid\bm{\theta})\mid\mathcal{D})$, and $\pi(p_{(u,v)}(s,t\mid\bm{\theta})\mid\mathcal{D})$, respectively.
@@ 3.3 Model uncertainty
Statistical literature is very rich in many and different proposals that address model uncertainty. We extend the proposal of García-Donato et al (2023) in Cox regression to competing risk models, and employ spike-and-slab methods for variable selection. Both procedures use a strategy whose first stage proposes a variable selection method and a subsequent analysis of the estimated model.
The approach proposed in García-Donato et al. (2023) is based on computing the posterior probability associated to each of the possible models (model space), from the null model containing no covariates to the full model with all covariates. The calculation of such posterior distribution uses a discrete uniform prior distribution over the model space and a multivariate normal distribution as the prior for the regression coefficients $\pi(\bm{\beta}_{k})=\mbox{N}(\bm{0},\,\Sigma),$ where $\Sigma^{-1}$ is expressed in terms of the precision matrix of the covariates. This variance-covariance matrix is based on the literature of conventional priors and the expected Fisher information matrix evaluated on the null model. Model search in the different $2^{p_{k}}$ models, where $p_{k}$ is the number of covariates in the cause-specific hazard for cause $k$, $k=1,\ldots,K$, is based on Gibbs sampling algorithm and the Rao-Blackwellized estimator (Gelfand and Smith, 1990; Ghosh and Clyde, 2011) is used for the estimation of the inclusion probability associated to each covariate.
The second selection method we will follow is based on spike and slab prior distributions for the set of regression coefficients. In this approach, the prior distribution associated to the $\bm{\beta}_{k}$ regression coefficients associated to cause $k$ is expressed as (Mitchell and Beauchamp, 1988)
$\pi(\bm{\beta}_{k}\mid\bm{\gamma})$ =$_{j=1}^{p_{k}}\,[(1-\gamma_{j})I_{0}(\beta_{kj})+\gamma_{j}\pi(\beta_{kj})],$
$\pi(\bm{\gamma}\mid\eta)$ =$_{j=1}^{p_{k}}\,\eta^{\gamma_{j}}(1-\eta)^{1-\gamma_{j}},$
@@
where $I_{0}(\beta_{kj})$ is a mass point at zero that accompanies the spike, $\pi(\beta_{kj})$ is a wide density, and $\bm{\gamma}$ is a $\{0,1\}$ vector which defines the different $2^{p_{k}}$ models.
@@
DIC measures the fit of the data to the model via the deviance with a penalty for the complexity of the model, the effective number of parameters. DIC has been a very popular procedure in the Bayesian world due to its practical relationship with MCMC procedures to approximate the posterior distribution and its good performance, basically in non-hierarchical scenarios as it is our case. Smaller values of DIC indicate a better-fitting model. WAIC criterion is an extension of the Akaike Information Criterion with a fully Bayesian conception and very good properties.
@@
$f(t\mid\mathcal{D}^{-(i)})=\int\,f(t\mid\bm{\theta})\,\pi(\bm{\theta}\mid\mathcal{D}^{-(i)})\,\mbox{d}\bm{\theta},$
where the $\mathcal{D}^{(-i)}$ denotes the data without the $i$th observation (leave-one-out procedure), and $f(t\mid\bm{\theta})$ is the subsequent probability sampling density. In the case of univariate survival times with right censoring, the CPO associated to observation $t_{i}$ is the value of that posterior marginal distribution evaluated at the observation $t_{i}$ if this time is observed and the posterior probability $P(T_{i}>t_{i}\mid\mathcal{D}^{(-i)})$ when $t_{i}$ is censored (Laud, 2014). Large CPOs values support the selected model because indicate a good tuning between the data and the model.
The above definition of CPO does not apply to competitive risk models in which there are different event causes and censoring. We propose an extension of the definition of the CPO associated to the $i$-th observation $t_{i}$ in a competing risk model as follows:
@@
    $P(T>t\mid\mathcal{D}^{-(i)})=S(t\mid\mathcal{D}^{-(i)})=\int S(t\mid\bm{\theta})\,\pi(\bm{\theta}\mid\mathcal{D}^{-(i)})\,\mbox{d}\bm{\theta},$
    where $S(t\mid\bm{\theta})$ is the overall survival function in(3).
@@
    $\ f_{k}(t\mid\mathcal{D}^{-(i)})=\int f_{k}(t\mid\bm{\theta})\,\pi(\bm{\theta}\mid\mathcal{D}^{-(i)})\,\mbox{d}\bm{\theta}=\int h_{k}(t\mid\bm{\theta})\,S(t\mid\bm{\theta})\,\pi(\bm{\theta}\mid\mathcal{D}^{-(i)})\,\mbox{d}\bm{\theta},$
    where $f_{k}(t\mid\bm{\theta})=h_{k}(t\mid\bm{\theta})\,S(t\mid\bm{\theta})$ is known as the subdensity sampling function for time-to-event $k$, defined in terms of the cause-specific hazard function for cause $k$ in (1) and the overall survival function (3).
CPOs and LPML ca be used to compare the predictive performance of different models as well as to search for outliers. We will use an approximation for the calculation of the CPOs based on the MCMC outputs (Ntzoufras, 2008).
4 Bayesian competing risk analysis in action: Cardiovascular and non-cardiovascular death

### v2

@@ Abstract
Competing risks models are survival models with several events of interest acting in competition and whose occurrence is only observed for the event that occurs first in time. This paper presents a Bayesian approach to these models in which the issue of model selection is treated in a special way by proposing generalizations of some of the Bayesian procedures used in univariate survival analysis. This research is motivated by a study on the survival of patients with heart failure undergoing cardiac resynchronization therapy, a procedure which involves the implant of a device to stabilize the heartbeat. Two different causes of death have been considered: cardiovascular and non-cardiovascular, and a set of baseline covariates are examined in order to better understand their relationship with both causes of death. Model selection procedures and model checking analyses have been implemented and assessed. The posterior distribution of some relevant outputs such as the overall survival function, cumulative incidence functions, and transition probabilities have been computed and discussed.
@@ 1 Introduction
The statistical framework for the aforementioned objectives is survival analysis. In particular, competing risks models in which the event of interest, in our case death, can occur from a variety of competing causes. The statistical literature is abundant regarding procedures for modeling competing survival scenarios (Putter et al., 2007; Chen et al., 2014). In fact, competing risks models have been widely applied to numerous settings, such treatment evaluation in clinical trials (Ameis et al., 2024), breast cancer (Basu and Tiwari, 2009) or hepatic diseases (Yu and Huang, 2023), among other relevant clinical topics.
In our case, we will focus on cause-specific risk modelling to quantitatively assess the behaviour of the survival process through the cause-specific hazard function for each cause of death, the overall survival function for the time in which death occurs, and the cumulative incidence function of that time due to each of the causes considered among others (Chen et al., 2014). We also incorporate a look at the modelling from the point of view of multistate models (Andersen et al., 2001). These are stochastic processes of continuous time and discrete state space representing the different health conditions considered. We discuss the transition probabilities between the different states and specially conditional transition probabilities when it is known that a patient remains alive a certain time after the CRT treatment. We will work within the Bayesian statistical methodology. This allows quantifying all sources of uncertainty associated with a statistical problem in a natural way through probability distributions.
The rest of the paper is organized as follows. Section 2 presents a retrospective follow-up cardiac resynchronization therapy study and a brief description of the subsequent dataset. Section 3 contains general information about the competing risks model, the main elements of the cause-specific risk approach, and the general Bayesian inferential process. Subsection 3.3, devoted to model uncertainty, presents a special variable selection scenario. In this topic, we extend the proposal by García García-Donato et al (2023) in Cox regression models for computing Bayes factors to competing risks models and employ spike-and-slab selection methods. In addition, we also expand the definition of conditional predictive ordinate (CPO) (Ibrahim et al., 2001) to the general competing risks framework which is needed in order to apply the logarithm of the pseudomarginal likelihood as a tool for model assessment. Section 4 returns to the study of CRT therapy and applies the procedures and proposals of Section 3 to the data of this problem. Finally, Section 5 includes some conclusions.
@@ 2 The cardiac resynchronization therapy study
Clinical follow-up was completed in all patients, from the time each individual underwent CRT implantation (time zero) until death, cardiovascular (CV) or non-cardiovascular (N-CV), or until the date of the end of the follow-up. Survival times were observed in the first case and were right-censored in the latter. The median and interquartile ranges of survival times were 2.75 and (1.2, 5.2) years, respectively. Survival times and times-to-event for each cause of death densities are represented in Figure 1.
Figure 1: Distributions of times-to-event. In (A) times to cardiovascular and non-cardiovascular deaths in red and blue, respectively; in (B) times to censoring.
Concrete cause of death was registered for each patient. CV death involves causes as stroke and heart ischemic attacks; while NCV deaths include different causes such as bleedings, cancer or infections. In all cases, the exact cause of death was effectively registered and grouped into CV or NCV.
@@
Table 1: Descriptive summary of study covariates for censored individuals, for those who died, from CV or N-CV causes, and for all of them together. In the case of quantitative covariates, the sample mean and standard deviation are presented; for categorical variables, the number of patients in the sample in each category and their percentage are reported.
@@
Previous AF
- Yes 83 (40.89%) 31 (56.36%) 24 (63.16%) 138 (46.62%)
- No 120 (59.11%) 24 (43.64%) 14 (36.84%) 158 (53.38%)
Sex
- Man 150 (73.89%) 48 (87.27%) 29 (76.32%) 227 (76.69%)
- Woman 53 (26.11%) 7 (12.73%) 9 (23.68%) 69 (23.31%)
Quantitative covariates.
Age 68.58 (10.00) 72.27 (7.73) 74.18 (7.15) 69.99 (9.51)
Creatinine 1.24 (0.45) 1.70 (0.84) 1.39 (0.49) 1.34 (0.57)
@@
The standard Kaplan-Meier procedures for describing survival probabilities do not provide proper results in the context of competing risks (Pepe and Mori, 1993). In these settings, an alternative measure is the cumulative incidence function for each cause, defined as the cumulative probability of failure from a specific cause before a certain time. Figure 2 presents the cumulative incidence function for CV and N-CV death computed from the R package cmprsk (Gray, 2022). Specifically, this cumulative incidence function was calculated as described in Gray, 1998. CV and N-CV death behave quite similarly, although cardiovascular death is always associated with a higher probability. It is important to note that after five years of the CRT intervention the difference between the two probabilities increases: between 0.15 and 0.20 for N-CV death and between 0.20-0.25 for CV death. From 7-8 years onwards there is an abrupt rise in CV death, while CV remains at around 0.2. The small number of patients we have from 10 years onwards does not produce consistent informative signals.
@@ 3 Bayesian competing risks modelling
3.1 Competing risks sampling model
Competing risks models are survival models in which the initiating event of the process contemplate different terminating events, which compete with each other to be the event that eventually occurs. There are different models suitable for analyzing competing risks problems, being the most used the multivariate time-to-events model (Gail, 1975), the sub-distribution model (Gray, 1988) or the cause-specific hazards model (Gaynor et al., 1993). The most popular, and the one we will use in our study, is the latter: the cause-specific hazards modelling. An updated review on this models and some applications can be found in Chen et al., 2014.
We assume $K$ terminating events or causes of failure. We define $T_{k}$ as the random variable that describes the time to the terminating event $k$, and $\delta$ as the indicator variable for the event that has occurred. In particular, $\delta=k$ when the event $k$ has occurred and $\delta=0$ indicating the situation in which no event of interest has been recorded during the study period and, therefore, we have right censored information. Let $T=\mbox{min}\{T_{1},\ldots,T_{K}\}$. Cause-specific hazard modelling uses as a basic concept the joint distribution of $T$ and $\delta$ through the so-called cause-specific hazard function, which represents the hazard of failing from a given event in the company of all other competing events. The cause-specific hazard function for $T$ due to cause $k$ is defined as
$h_{k}(t\mid\boldsymbol{\theta}_{k})=\lim_{\varDelta t\to 0}\frac{P(t\leq T<t+\varDelta t,\delta=k\mid T\geq t,\,\boldsymbol{\theta}_{k})}{\varDelta t},$ (1)
where $\boldsymbol{\theta}_{k}$ is the subsequent parametric vector. It is important to remember that our inferential framework is Bayesian statistics. Consequently, the uncertainty associated with the parameters will be expressed in probabilistic terms and for this reason we will express in terms of conditional probability distributions all probabilistic elements, such as $h_{k}(t\mid\boldsymbol{\theta_{k}})$, that depend on parameters.
The most common modeling of the cause-specific hazard function for cause $k$ is through a proportional hazards model as follows
$h_{k}(t\mid\boldsymbol{\theta}_{k})=h_{0,k}(t\mid\boldsymbol{\phi}_{k})\>\exp\{\boldsymbol{x}^{\top}_{k}\,\boldsymbol{\beta}_{k}\},$ (2)
where $h_{0,k}(t\mid\boldsymbol{\phi_{k}})$ is a baseline hazard function that depends on time and a parametric vector $\boldsymbol{\phi}_{k}$. It expresses the shape of the cause-specific hazard function. The vector $\boldsymbol{\beta}_{k}$ represents the regression coefficients associated to covariates $\boldsymbol{x}_{k}$ of covariates, and $\boldsymbol{\theta}_{k}=(\boldsymbol{\phi}_{k},\boldsymbol{\beta}_{k})$. Baseline hazard functions can be established through parametric or semiparametric approaches. Parametric models generate functions with rigid shapes that cannot admit irregular behavior. On the other hand, semi-parametric modeling provides very flexible functions but suffer from overfitting and instability (Ibrahim et al., 2014).
@@ 3.1 Competing risks sampling model
where now $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{K})$.
@@
$F_{k}(t\mid\boldsymbol{\theta})=P(T\leq t,\>\delta=k\mid\boldsymbol{\theta})=\int_{0}^{t}h_{k}(u\mid\boldsymbol{\theta_{k}})\>S(u\mid\boldsymbol{\theta})\>\mbox{d}u.$
This function provides the probability associated with the occurrence of each of the competing events, $P(\delta=k\mid\boldsymbol{\theta})=\lim_{t\to\infty}F_{k}(t\mid\boldsymbol{\theta})$.
Competing risks models can also be seen as a particular multi-state model. In general, a multi-state process is a continuous time stochastic process $\{Z(t),t\geq 0\}$ with a finite state space $\Omega=\{0,1,\ldots,K\}$, where $Z(t)=k$ indicates that the process is in state $k$ at time $t$, or in terms of survival analysis that event $k$ has occurred at some time instant less than $t$. The initial distribution of the process is $P(Z(0)=0)=1$ and relevant transition probabilities between states are
$p_{(0,0)}(0,t\mid\boldsymbol{\theta})$ $=P(Z(t)=0\mid Z(0)=0,\boldsymbol{\theta})=P(T>t\mid\boldsymbol{\theta})=S(t\mid\boldsymbol{\theta}),$
$p_{(0,k)}(0,t\mid\boldsymbol{\theta})$ $=P(Z(t)=k\mid Z(0)=0,\boldsymbol{\theta})=P(T\leq t,\delta=k\mid\boldsymbol{\theta})=F_{k}(t\mid\boldsymbol{\theta}),\,\,k\neq 0.$
@@
$p_{(0,0)}(s,t\mid\boldsymbol{\theta})=$ $\,\frac{P(Z(t)=0,Z(s)=0\mid\boldsymbol{\theta})}{P(Z(s)=0\mid\boldsymbol{\theta})}=\frac{S(t\mid\boldsymbol{\theta})}{S(s\mid\boldsymbol{\theta})},\,t\geq s,$
$p_{(0,k)}(s,t\mid\theta)=$ $\,\frac{P(Z(t)=k,Z(s)=0\mid\boldsymbol{\theta})}{P(Z(s)=0\mid\boldsymbol{\theta})}=\frac{P(s<T<t,\delta=k\mid\boldsymbol{\theta})}{S(s\mid\boldsymbol{\theta})}$
$=$ $\,\frac{F_{k}(t\mid\boldsymbol{\theta})-F_{k}(s\mid\boldsymbol{\theta})}{S(s\mid\boldsymbol{\theta})},\,t\geq s,\,k\neq 0.$
@@ 3.2 Prior and posterior distributions
Bayesian models include a sampling model and a prior distribution $\pi(\boldsymbol{\theta})$ for the unknown parameters that describes how much, little or no information is available about $\boldsymbol{\theta}$. We will work in a general scenario of prior independence between the parameters associated with the different competing risks as well as between the parameters governing the time-dependent structure of each hazard function and the regression coefficients of the variables, $\pi(\boldsymbol{\theta})=\prod_{k=1}^{K}\,\pi(\boldsymbol{\phi}_{k},\boldsymbol{\beta}_{k})=\prod_{k=1}^{K}\,\pi(\boldsymbol{\phi}_{k})\,\pi(\boldsymbol{\beta}_{k})$.
The database is built on the basis that survival times due to one of the causes provides censored information on the survival times of the other competing causes. The information provided by the data $\mathcal{D}$ is described in terms of the likelihood function of $\boldsymbol{\theta}$, $L(\boldsymbol{\theta})$, which is calculated through the hazard function (Chen et al., 2014). The prior distribution and the likelihood function are combined to obtain the posterior distribution of $\boldsymbol{\theta}$, $\pi(\boldsymbol{\theta}\mid\mathcal{D})$. This distribution is not analytical but an approximate sample $\{\boldsymbol{\theta}^{(1)},\ldots,\boldsymbol{\theta}^{(M)}\}$ of such a distribution can be obtained through Markov chain Monte Carlo, MCMC, methods. This distribution includes all available information and is the starting point for the analysis of the problem.
The overall survival function, cumulative incidence function, probability associated with the occurrence of each competing event as well as transitions probabilities presented before are relevant outputs that depend on $\boldsymbol{\theta}$. Consequently, we can use the posterior approximate sample from $\pi(\boldsymbol{\theta}\mid\mathcal{D})$ to generate approximate random samples for their subsequent posteriors, $\pi(S(t\mid\boldsymbol{\theta})\mid\mathcal{D})$, $\pi(F_{k}(t\mid\boldsymbol{\theta})\mid\mathcal{D})$, and $\pi(p_{(u,v)}(s,t\mid\boldsymbol{\theta})\mid\mathcal{D})$, respectively.
@@ 3.3 Model uncertainty
Statistical literature is very rich in many and different proposals that address model uncertainty. We extend the proposal of García-Donato et al (2023) in Cox regression to competing risks models, and employ spike-and-slab methods for variable selection. Both procedures use a strategy whose first stage proposes a variable selection method and a subsequent analysis of the estimated model.
The approach proposed in García-Donato et al. (2023) is based on computing the posterior probability associated to each of the possible models (model space), from the null model containing no covariates to the full model with all covariates. The calculation of such posterior distribution uses a discrete uniform prior distribution over the model space and a multivariate normal distribution as the prior for the regression coefficients $\pi(\boldsymbol{\beta}_{k})=\mbox{N}(\boldsymbol{0},\,\Sigma),$ where $\Sigma^{-1}$ is expressed in terms of the precision matrix of the covariates. This variance-covariance matrix is based on the literature of conventional priors and the expected Fisher information matrix evaluated on the null model. Model search in the different $2^{p_{k}}$ models, where $p_{k}$ is the number of covariates in the cause-specific hazard for cause $k$, $k=1,\ldots,K$, is based on Gibbs sampling algorithm and the Rao-Blackwellized estimator (Gelfand and Smith, 1990; Ghosh and Clyde, 2011) is used for the estimation of the inclusion probability associated to each covariate.
The second selection method we will follow is based on spike and slab prior distributions for the set of regression coefficients. In this approach, the prior distribution associated to the $\boldsymbol{\beta}_{k}$ regression coefficients associated to cause $k$ is expressed as (Mitchell and Beauchamp, 1988)
$\pi(\boldsymbol{\beta}_{k}\mid\boldsymbol{\gamma})$ =$_{j=1}^{p_{k}}\,[(1-\gamma_{j})I_{0}(\beta_{kj})+\gamma_{j}\pi(\beta_{kj})],$
$\pi(\boldsymbol{\gamma}\mid\eta)$ =$_{j=1}^{p_{k}}\,\eta^{\gamma_{j}}(1-\eta)^{1-\gamma_{j}},$
@@
where $I_{0}(\beta_{kj})$ is a mass point at zero that accompanies the spike, $\pi(\beta_{kj})$ is a wide density, and $\boldsymbol{\gamma}$ is a $\{0,1\}$ vector which defines the different $2^{p_{k}}$ models.
@@
DIC measures the fit of the data to the model via the deviance with a penalty for the complexity of the model, the effective number of parameters. This criterion has been a very popular procedure in the Bayesian world due to its practical relationship with MCMC procedures to approximate the posterior distribution and its good performance, basically in non-hierarchical scenarios as it is our case. Smaller values of DIC indicate a better-fitting model. DIC is defined as $\text{DIC}=D(\bar{\theta})+2pD$, where $D(\theta)$ is the deviance of the posterior distribution of $\theta$, and $pD$ is the effective number of parameters and it is calculated as $pD=\text{E}[D(\theta)]-D(\bar{\theta})$.
WAIC criterion is an extension of the Akaike Information Criterion with a fully Bayesian conception. Its definition is similar to the DIC, as it assesses the model fit and penalizes models with a high number of parameters. WAIC has been calculated as $\text{WAIC}=-2\sum_{i=1}^{n}\log p_{post(y_{i})}+2pD$, where $\sum_{i=1}^{n}\log p_{post(y_{i})}$ is the sum of predictive density for each data point and $pD$ the number of effective parameters. Although in the context of non-hierarchical modeling like this one DIC and WAIC should provide similar information, we have obtained both indexes for the sake of completeness. As expected, they have always reached similar conclusions.
@@
$f(t\mid\mathcal{D}^{-(i)})=\int\,f(t\mid\boldsymbol{\theta})\,\pi(\boldsymbol{\theta}\mid\mathcal{D}^{-(i)})\,\mbox{d}\boldsymbol{\theta},$
where the $\mathcal{D}^{(-i)}$ denotes the data without the $i$th observation (leave-one-out procedure), and $f(t\mid\boldsymbol{\theta})$ is the subsequent probability sampling density. In the case of univariate survival times with right censoring, the CPO associated to observation $t_{i}$ is the value of that posterior marginal distribution evaluated at the observation $t_{i}$ if this time is observed and the posterior probability $P(T_{i}>t_{i}\mid\mathcal{D}^{(-i)})$ when $t_{i}$ is censored (Laud, 2014). Large CPOs values support the selected model because indicate a good tuning between the data and the model.
The above definition of CPO does not apply to competitive risk models in which there are different event causes and censoring. We propose an extension of the definition of the CPO associated to the $i$-th observation $t_{i}$ in a competing risks model as follows:
@@
    $P(T>t\mid\mathcal{D}^{-(i)})=S(t\mid\mathcal{D}^{-(i)})=\int S(t\mid\boldsymbol{\theta})\,\pi(\boldsymbol{\theta}\mid\mathcal{D}^{-(i)})\,\mbox{d}\boldsymbol{\theta},$
    where $S(t\mid\boldsymbol{\theta})$ is the overall survival function in(3).
@@
    $\ f_{k}(t\mid\mathcal{D}^{-(i)})=\int f_{k}(t\mid\boldsymbol{\theta})\,\pi(\boldsymbol{\theta}\mid\mathcal{D}^{-(i)})\,\mbox{d}\boldsymbol{\theta}=\int h_{k}(t\mid\boldsymbol{\theta})\,S(t\mid\boldsymbol{\theta})\,\pi(\boldsymbol{\theta}\mid\mathcal{D}^{-(i)})\,\mbox{d}\boldsymbol{\theta},$
    where $f_{k}(t\mid\boldsymbol{\theta})=h_{k}(t\mid\boldsymbol{\theta})\,S(t\mid\boldsymbol{\theta})$ is known as the subdensity sampling function for time-to-event $k$, defined in terms of the cause-specific hazard function for cause $k$ in (1) and the overall survival function (3).
CPOs and LPML can be used to compare the predictive performance of different models as well as to search for outliers. We will use an approximation for the calculation of the CPOs based on the MCMC outputs (Ntzoufras, 2008).
4 Bayesian competing risks analysis in action: Cardiovascular and non-cardiovascular death